OpenStudy (anonymous):

find the product of (-8)x(6)x(-2)

OpenStudy (tkhunny):

Do you mean \(-8x^{6}x^{-2}\)?

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ (-8) \times (6) \times (-2) }\) like this ?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

multiply them just like you were to multiply together 2,6 and 8. The sign will be positive, since there are even negatives in the product.

OpenStudy (anonymous):

144?

OpenStudy (tkhunny):

Why are you guessing? 8*6*2 = 8*12 = ?

OpenStudy (anonymous):

96

OpenStudy (anonymous):

ok got it thx